# Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty

### v1

@@
Rethinking Arrow–Debreu: A New Framework for Exchange, Time, and Uncertainty
Nizar Riane
@@ 1 Introduction
The Arrow–Debreu Model (ADM) [AD54] has long been a cornerstone of modern economic theory, providing a rigorous mathematical framework for analyzing microeconomic decisions. It offers a formal resolution to an old question posed by Léon Walras in [Wal74], leveraging advancements in the theory of multifunctions and fixed points.
@@
While the ADM is meticulously formulated, it has not been universally embraced by economists. Criticism often targets its assumptions, including perfect competition, the existence of forward markets for every commodity and all conceivable contingencies, and the absence of money as a store of value. This has led some scholars, such as Mark Blaug, to express strong reservations. Blaug famously characterized Arrow and Debreu’s seminal paper [AD54] as marking "the beginning of what has since become a cancerous growth in the very center of microeconomics" (see [Bla98]). He further argued that "by the time we got to Arrow and Debreu, general equilibrium theory had ceased to make any descriptive claim about actual economic systems and had become a purely formal apparatus about a quasi-economy". His criticism extended to Gérard Debreu’s book, The Theory of Value [Deb59], which he deemed "probably the most arid and pointless book in the entire literature of economics".
@@
Beyond Blaug’s critiques of its theoretical abstraction, it is important to recognize that general equilibrium theory has undergone significant refinements since the publication of Arrow and Debreu’s 1954 paper. Debreu’s The Theory of Value [Deb59] further formalized the framework, particularly in economies facing uncertainty. The role of money within the model has been extensively debated, with insightful contributions from Robert Clower [Clo67], Frank Hahn [Hah89], and Joseph M. Ostroy and Ross M. Starr [OS74] and others. The question of equilibrium convergence was explored using nontraditional techniques and more realistic approaches (see, for example, [HN62]). Additionally, time was explicitly incorporated into the system [Rad72], and market imperfections-such as equilibrium with indivisible goods-were examined to better align the theory with economic realities [Sve84].
@@
In his book The Methodology of Economics: Or, How Economists Explain [Bla92], Mark Blaug provocatively asks whether general equilibrium theory should be regarded as an empirical theory or merely as a conceptual framework. To explore this question, one can draw an analogy with classical physics: just as Newtonian mechanics assumes the existence of a Galilean reference frame–a perfectly inertial system in which the laws of motion take their simplest form–general equilibrium theory is grounded in idealized assumptions of complete markets and perfect competition. However, in practice, the Earth is a non-inertial frame of reference: its rotation gives rise to apparent forces, such as the Coriolis effect, which must be accounted for to correctly describe physical phenomena. Likewise, real-world markets are subject to imperfections and incompleteness, forming what one might call a "non-Galilean" economic environment. In such contexts, the neat predictions of general equilibrium theory must be modified to incorporate frictions, constraints, and institutional complexities. Thus, while general equilibrium provides a powerful and elegant framework for understanding economic coordination, its empirical relevance depends on our ability to account for the deviations from its foundational assumptions.
@@
Utility in real-world exchanges depends on effective allocations-what individuals receive or consume through transactions-rather than theoretical allocations. This distinction underscores the importance of addressing the outcomes of exchanges rather than unattainable aspirations. Markets are not arenas for idealized redistribution but mechanisms for facilitating feasible transactions shaped by constraints and negotiation.
@@
Moreover, the quantity exchanged in a transaction is inherently constrained. It is determined by the minimum that both parties can support, reflecting their respective limitations. This contrasts with theoretical models that emphasize optimal allocations without considering such practical constraints.
@@
The Rational Expectations Hypothesis, which assumes that agents form beliefs based on the true statistical expectation of future variables, has long served as a foundational assumption in anticipation modeling. However, this framework abstracts away from key psychological and informational constraints faced by real-world decision-makers. In practice, individuals do not bet on the future by averaging over full probability distributions. Instead, their anticipations are shaped by bounded rationality, limited attention, memory constraints, and heuristic reasoning. And one should distinguish between anticipation an uncertain future and gambling over a set of lotteries. In light of these limitations, we propose a departure from the traditional use of expectations and instead define anticipation as the conditional mode—that is, the most likely realization of a variable given the agent’s information, memory, and subjective beliefs. This approach better reflects the role of instinctive judgment, focal outcomes, and perceived salience, aligning with Keynes’ notion of "animal spirits" [Key36]. By grounding anticipations in the mode rather than the mean, this formulation introduces behavioral realism without sacrificing formal rigor. It allows for agent heterogeneity, evolving beliefs, and more nuanced responses to uncertainty—features that are essential in dynamic, strategic environments.
@@ 2 New way to consider the Consumer Problem
A strict inequality would imply that the customer accepted an exchange in which higher-valued goods were traded for lower-valued ones, which effectively equivalent to setting a price lower than $p^{i}$. One might consider a condition similar to that in the Arrow–Debreu model [AD54]:
@@
Definition of a Generalized Game[Kre12]
@@
Equilibrium in Effective Trade Economy[Kre12]
@@
Pareto Optimal Allocation [Sta12]
@@ 3 KKT characterization
In this section we make the following assumption:
@@
And for simplicity, let $\mathscr{X}^{i}={\mathbb{R}}_{+}^{L}$ and denote $x^{i}$ as $x$ and $x^{j}$ as $y$, respectively. Define the Lagrangian function $\mathcal{L}(p^{i},X,\lambda,\mu,\eta,\nu,\theta,\kappa)$ as
@@
$\frac{\partial\mathcal{L}}{\partial p^{i}_{k}}$ $=\lambda\sum_{j\neq i}x_{ij,k}+\nu_{k}+\kappa=0$
$\frac{\partial\mathcal{L}}{\partial x_{ij,k}}$ $=-\frac{\partial u}{\partial x_{k}}+\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}=0$
$\frac{\partial\mathcal{L}}{\partial x_{ji,k}}$ $=\frac{\partial u}{\partial x_{k}}-\lambda p^{j}_{k}+\mu_{ji,k}-\eta_{ji,k}+\theta_{k}=0\,.$
@@
$\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa$
$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}>0$
$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{j}_{k}-\mu_{ji,k}+\eta_{ji,k}-\theta_{k}>0\,.$
This implies, for $i\neq j\neq l$:
@@
$\mu_{ij,k}-\eta_{ij,k}$ $=\mu_{il,k}-\eta_{il,k}$
$\eta_{ji,k}-\mu_{ji,k}$ $=\eta_{li,k}-\mu_{li,k}\,.$

### v2

@@
Abstract
This paper revisits the Arrow-Debreu general equilibrium framework through the lens of effective trade, emphasizing the distinction between theoretical and realizable market interactions. We develop the Effective Trade Model (ETM), where transactions arise from bilateral feasibility rather than aggregate supply and demand desires. Within this framework, we establish the main properties of the price-demand correspondence and prove the existence of Nash equilibria, incorporating production, money, and network topology. The analysis extends to time, uncertainty, and open economies, revealing how loanable funds and exchange rates emerge endogenously. Our results show that equilibrium is shaped by transaction constraints, subjective pricing, and decentralized negotiation, rather than by universal market-clearing conditions, and thereby call into question the foundations of welfare theory. Anticipation is modeled via the conditional mode, capturing bounded rationality and information limitations in contrast to the rational expectations hypothesis. The ETM thus offers a behaviorally and structurally grounded alternative to classical general equilibrium, bridging microfoundations, monetary dynamics, and temporal consistency within a unified framework.
@@ 1 Introduction
The Arrow–Debreu Model (ADM) [Arrow and Debreu, 1954] has long been a cornerstone of modern economic theory, providing a rigorous mathematical framework for analyzing microeconomic decisions. It offers a formal resolution to an old question posed by Léon Walras in [Walras, 1874], leveraging advancements in the theory of multifunctions and fixed points.
@@
While the ADM is meticulously formulated, it has not been universally embraced by economists. Criticism often targets its assumptions, including perfect competition, the existence of forward markets for every commodity and all conceivable contingencies, and the absence of money as a store of value. This has led some scholars, such as Mark Blaug, to express strong reservations. Blaug famously characterized Arrow and Debreu’s seminal paper [Arrow and Debreu, 1954] as marking "the beginning of what has since become a cancerous growth in the very center of microeconomics" (see [Blaug, 1998]). He further argued that "by the time we got to Arrow and Debreu, general equilibrium theory had ceased to make any descriptive claim about actual economic systems and had become a purely formal apparatus about a quasi-economy". His criticism extended to Gérard Debreu’s book, The Theory of Value [Debreu, 1959], which he deemed "probably the most arid and pointless book in the entire literature of economics".
@@
Beyond Blaug’s critiques of its theoretical abstraction, it is important to recognize that general equilibrium theory has undergone significant refinements since the publication of Arrow and Debreu’s 1954 paper. Debreu’s The Theory of Value [Debreu, 1959] further formalized the framework, particularly in economies facing uncertainty. The role of money within the model has been extensively debated, with insightful contributions from Robert Clower [Clower, 1967], Frank Hahn [Hahn, 1989], and Joseph M. Ostroy and Ross M. Starr [Ostroy and Starr, 1974] and others. The conceptualization of transaction was reformulated by L. Shapley and M. Shubik through the introduction of trading posts in [ShapleyShubik1977]. The question of equilibrium convergence was explored using nontraditional techniques and more realistic approaches (see, for example, [Hahn and Negishi, 1962]). Additionally, time was explicitly incorporated into the system [Radner, 1972], and market imperfections-such as equilibrium with indivisible goods-were examined to better align the theory with economic realities [Svensson, 1984].
@@
In his book The Methodology of Economics: Or, How Economists Explain [Blaug, 1992], Mark Blaug provocatively asks whether general equilibrium theory should be regarded as an empirical theory or merely as a conceptual framework. To explore this question, one can draw an analogy with classical physics: just as Newtonian mechanics assumes the existence of a Galilean reference frame–a perfectly inertial system in which the laws of motion take their simplest form–general equilibrium theory is grounded in idealized assumptions of complete markets and perfect competition. However, in practice, the Earth is a non-inertial frame of reference: its rotation gives rise to apparent forces, such as the Coriolis effect, which must be accounted for to correctly describe physical phenomena. Likewise, real-world markets are subject to imperfections and incompleteness, forming what one might call a "non-Galilean" economic environment. In such contexts, the neat predictions of general equilibrium theory must be modified to incorporate frictions, constraints, and institutional complexities. Thus, while general equilibrium provides a powerful and elegant framework for understanding economic coordination, its empirical relevance depends on our ability to account for the deviations from its foundational assumptions.
@@
Utility in real-world exchanges depends on effective allocations-what individuals receive or consume through transactions-rather than theoretical allocations. This distinction underscores the importance of addressing the outcomes of exchanges rather than unattainable aspirations. Markets are not arenas for idealized redistribution but mechanisms for facilitating feasible transactions shaped by constraints, negotiation, and network interactions.
@@
Moreover, the quantity exchanged in a transaction is inherently constrained. It is determined by the minimum of what both parties can support, reflecting their respective limitations. This contrasts with theoretical models that emphasize optimal allocations without considering such practical constraints.
@@
The Rational Expectations Hypothesis, which assumes that agents form beliefs based on the true statistical expectation of future variables, has long served as a foundational assumption in anticipation modeling. However, this framework abstracts away from key psychological and informational constraints faced by real-world decision-makers. In practice, individuals do not bet on the future by averaging over full probability distributions. Instead, their anticipations are shaped by bounded rationality, limited attention, memory constraints, and heuristic reasoning. And one should distinguish between anticipation an uncertain future and gambling over a set of lotteries. In light of these limitations, we propose a departure from the traditional use of expectations and instead define anticipation as the conditional mode—that is, the most likely realization of a variable given the agent’s information, memory, and subjective beliefs. This approach better reflects the role of instinctive judgment, focal outcomes, and perceived salience, aligning with Keynes’ notion of "animal spirits" [Keynes, 1936]. By grounding anticipations in the mode rather than the mean, this formulation introduces behavioral realism without sacrificing formal rigor. It allows for agent heterogeneity, evolving beliefs, and more nuanced responses to uncertainty—features that are essential in dynamic, strategic environments.
@@ 2 New way to consider the Consumer Problem
A strict inequality would imply that the customer accepted an exchange in which higher-valued goods were traded for lower-valued ones, which effectively equivalent to setting a price lower than $p^{i}$. One might consider a condition similar to that in the Arrow–Debreu model [Arrow and Debreu, 1954]:
@@
Definition of a Generalized Game[Kreps, 2012]
@@
Equilibrium in Effective Trade Economy[Kreps, 2012]
@@
Pareto Optimal Allocation [Starr, 2012]
@@ 3 KKT characterization
We now characterize the equilibrium for a smooth economy, that is, when the relevant functions are well-behaved. We make the following assumption:
@@
And for simplicity, let $\mathscr{X}^{i}={\mathbb{R}}_{+}^{L}$ and denote $x^{i}$ as $x$ and $x^{j}$ as differentiable function $y(p^{i},X^{i})$ of $p^{i}$ and $X^{i}$, respectively. Define the Lagrangian function $\mathcal{L}(p^{i},X,\lambda,\mu,\eta,\nu,\theta,\kappa)$ as
@@
$\frac{\partial\mathcal{L}}{\partial p^{i}_{k}}$ $=\lambda\sum_{j\neq i}x_{ij,k}+\nu_{k}+\kappa+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial p^{i}_{k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial p^{i}_{k}}y_{ji,l}=0$
$\frac{\partial\mathcal{L}}{\partial x_{ij,k}}$ $=-\frac{\partial u}{\partial x_{k}}+\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ij,k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ij,k}}y_{ji,l}=0$
$\frac{\partial\mathcal{L}}{\partial x_{ji,k}}$ $=\frac{\partial u}{\partial x_{k}}-\lambda p^{j}_{k}+\mu_{ji,k}-\eta_{ji,k}+\theta_{k}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ji,k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ji,k}}y_{ji,l}=0\,.$
@@
$\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial p^{i}_{k}}y_{ij,l}-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial p^{i}_{k}}y_{ji,l}$
$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{i}_{k}+\mu_{ij,k}-\eta_{ij,k}-\theta_{k}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ij,k}}y_{ij,l}+\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ij,k}}y_{ji,l}$
$\frac{\partial u}{\partial x_{k}}$ $=\lambda p^{j}_{k}-\mu_{ji,k}+\eta_{ji,k}-\theta_{k}-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial x_{ji,k}}y_{ij,l}-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial x_{ji,k}}y_{ji,l}\,.$
  1. 1.
    This implies, for $i\neq j$, the price-quantity equation:
@@
    $-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ij,l}+\frac{\partial}{\partial x_{ji,k}}y_{ij,l}\right)-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ji,l}+\frac{\partial}{\partial x_{ji,k}}y_{ji,l}\right).$
    Interpretation:
    3.0.0.1 Shadow valuation of price differences:
    The difference $p^{i}_{k}-p^{j}_{k}$ measures the relative valuation of good $k$ between the two trading partners. The multiplier $\lambda$ is the shadow value of consumer $i$’s trade-balance constraint
    $\sum_{k=1}^{L}p^{i}_{k}\sum_{j\neq i}x_{ij,k}-\sum_{k=1}^{L}\sum_{j\neq i}p^{j}_{k}x_{ji,k}$
    Thus, $\lambda(p^{i}_{k}-p^{j}_{k})$ quantifies the utility-valued shadow advantage or disadvantage that arises when consumer $i$ reallocates one marginal unit of good $k$ between “selling at his own price” and “selling at $j$’s price,” ceteris paribus. In short, How much does $i$ benefit (in utility) from trading good $k$ at his own price rather than $j$’s price.
    3.0.0.2 Bounds net direct marginal shadow price.
    The multipliers $\eta_{ij,k}$ and $\eta_{ji,k}$ are the shadow prices of the upper–bound constraints $x_{ij,k}\leq y_{ij,k}$ and $x_{ji,k}\leq y_{ji,k}$ imposed by consumer $j$. A positive $\eta_{ij,k}$ (resp. $\eta_{ji,k}$) therefore measures the marginal utility gain for consumer $i$ from relaxing $j$’s cap on how much he is willing to buy (resp. sell). The multipliers $\mu_{ij,k}$ and $\mu_{ji,k}$ are the shadow prices of the lower–bound constraints $x_{ij,k}\geq 0$ and $x_{ji,k}\geq 0$; they are positive exactly when $i$ is forced to remain at a zero–flow corner and would benefit, at the margin, from allowing a flow in the opposite direction.
    Thus
    $\eta_{ij,k}+\eta_{ji,k}-(\mu_{ij,k}+\mu_{ji,k})$
    is the net direct marginal shadow value of all bilateral quantity bounds on good $k$ in the transaction with $j$. It summarizes whether, at the margin, $i$’s welfare is primarily constrained by $j$’s upper trading limits (reflected in the $\eta$’s) or by the zero–flow lower bounds (reflected in the $\mu$’s), thereby capturing the overall tightness of the bilateral quantity restrictions faced by $i$.
    3.0.0.3 Network indirect adjustment costs:
    The quantities $y_{ij,\ell}$ and $y_{ji,\ell}$ denote consumer $j$’s own proposed trades. The derivatives $\frac{\partial y_{ij,\ell}}{\partial x_{ji,k}}$ and $\frac{\partial y_{ji,\ell}}{\partial x_{ji,k}}$ (resp. $\frac{\partial y_{ij,\ell}}{\partial x_{ij,k}}$ and $\frac{\partial y_{ji,\ell}}{\partial x_{ij,k}}$) measure how $j$’s proposals react when $i$ changes $x_{ji,k}$ (resp. $x_{ij,k}$). The sums
    $-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ij,l}+\frac{\partial}{\partial x_{ji,k}}y_{ij,l}\right)-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ji,l}\left(\frac{\partial}{\partial x_{ji,k}}y_{ji,l}+\frac{\partial}{\partial x_{ji,k}}y_{ji,l}\right)$
    capture the indirect marginal constraint costs that arise because altering $i$’s intended transactions of good $k$ affects the feasibility of other potential transactions through $j$’s strategic adjustments. Economically, these terms represent transaction-network externalities: changing one bilateral flow modifies the entire structure of feasible exchanges.
    3.0.0.4 Economic meaning of the equation:
    The KKT condition can be summarized as the identity
[…]
    Economically, this states that at an optimum consumer $i$ adjusts the bilateral flow of good $k$ with $j$ until the utility-valued marginal gain from reallocating one unit of the good across the price gap $p^{i}_{k}-p^{j}_{k}$ is exactly balanced by (i) the net shadow value of the binding quantity bounds in the $(i,j)$ relationship and (ii) the indirect constraint costs generated by how $j$’s adjustment of other proposed trades reacts to $i$’s change in this particular flow.
    In short, the marginal price advantage of trading with $j$ at $i$’s valuation rather than $j$’s valuation is driven to equality with the total marginal feasibility cost created by bilateral quantity limits and by the propagation of this adjustment through the transaction network.
  2. 2.
    Let’s analyze the price equation:
    $\lambda\sum_{j\neq i}x_{ij,k}$ $=-\nu_{k}-\kappa-\sum_{l=1}^{L}\sum_{j\neq i}\eta_{ij,l}\frac{\partial}{\partial p^{i}_{k}}y_{ij,l}-\sum_{k=1}^{L}\sum_{j\neq i}\eta_{ji,l}\frac{\partial}{\partial p^{i}_{k}}y_{ji,l}\geqslant 0$
    Interpretation:
    3.0.0.5 Shadow impact of the price of good $k$.
    The multiplier $\lambda$ is the shadow value of consumer $i$’s trade-balance constraint, so it converts any marginal perturbation of this constraint into utility units. The coefficient $\sum_{j\neq i}x_{ij,k}$ is the sensitivity of the trade-balance constraint to changes in the price component $p^{i}_{k}$. Thus the product $\lambda\sum_{j\neq i}x_{ij,k}$ represents the utility-valued marginal effect of increasing $p^{i}_{k}$ through its impact on consumer $i$’s trade-balance constraint. It is the utility-weighted exposure of the trade-balance condition to this price component.
    3.0.0.6 Nonnegativity constraint cost.
    The term $\nu_{k}$ is associated with the lower-bound constraint $p^{i}_{k}\geq 0$. If $p^{i}_{k}=0$, then $\nu_{k}>0$ indicates that consumer $i$ would strictly prefer to lower $p^{i}_{k}$ further, but the constraint prevents such a move. Hence $\nu_{k}$ represents the marginal utility loss created by the infeasibility of decreasing $p^{i}_{k}$ below zero; the negative sign reflects that this is a constraint-induced cost from the perspective of stationarity.
    3.0.0.7 Normalization constraint cost.
    The multiplier $\kappa$ enforces the unit-sum constraint $\sum_{r=1}^{L}p^{i}_{r}=1$. Raising $p^{i}_{k}$ necessarily requires reducing the weight placed on other prices. Thus
    $\kappa$
    captures the marginal shadow cost of transferring one unit of price weight toward $p^{i}_{k}$ while preserving the normalization constraint. It measures the opportunity cost of reallocating price mass across goods.
    3.0.0.8 Network reaction costs from partners’ demand responses.
    Consider the term
    $\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ij,\ell}\frac{\partial y_{ij,\ell}}{\partial p^{i}_{k}}.$
    The quantity $y_{ij,\ell}$ is consumer $j$’s proposed purchase of good $\ell$ from $i$. If increasing $p^{i}_{k}$ makes $j$ reduce $y_{ij,\ell}$, then the constraint $x_{ij,\ell}\leq y_{ij,\ell}$ becomes tighter. The multiplier $\eta_{ij,\ell}$ is the shadow value of that constraint. Thus this sum captures the marginal utility loss generated by the tightening of all purchase-side feasibility constraints caused by partners reducing their intended demands in response to a higher price $p^{i}_{k}$.
    3.0.0.9 Network reaction costs from partners’ supply responses.
    Similarly,
    $\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ji,\ell}\frac{\partial y_{ji,\ell}}{\partial p^{i}_{k}}$
    reflects how raising $p^{i}_{k}$ affects $j$’s proposed sales of good $\ell$ to $i$. If these sales decrease, then the constraint $x_{ji,\ell}\leq y_{ji,\ell}$ becomes more binding. Weighted by the multipliers $\eta_{ji,\ell}$, this term represents the corresponding marginal utility loss due to the tightening of supply-side feasibility constraints.
    3.0.0.10 Economic meaning of the full KKT condition.
    The stationarity condition
    $\lambda\sum_{j\neq i}x_{ij,k}=-\nu_{k}-\kappa-\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ij,\ell}\frac{\partial y_{ij,\ell}}{\partial p^{i}_{k}}-\sum_{\ell=1}^{L}\sum_{j\neq i}\eta_{ji,\ell}\frac{\partial y_{ji,\ell}}{\partial p^{i}_{k}}$
    states that consumer $i$ sets the price component $p^{i}_{k}$ so that the utility-valued marginal effect of altering $p^{i}_{k}$ through the trade-balance constraint is exactly offset by the aggregate marginal shadow costs that this change induces. These costs consist of: (i) the marginal cost of respecting the nonnegativity constraint on prices, (ii) the marginal opportunity cost induced by the normalization constraint, and (iii) the propagated tightening of all bilateral feasibility constraints caused by trading partners’ strategic adjustments to $i$’s price change.
    In equilibrium, $p^{i}_{k}$ is therefore chosen such that the marginal benefit of adjusting this price is precisely counterbalanced by the total marginal feasibility costs transmitted through the entire trading network.
4 Convergence
Next, we propose three convergence processes toward Pareto-optimal and Nash–Pareto states, based on two underlying principles: